OpenStudy (anonymous):

A circle has a diameter of 6 millimeters. What are the circumference and area of the circle? Use three point one four for pi. Round to the nearest tenth. Answer eighteen point eight millimeters; twenty-eight point three square millimeters nine point four millimeters; twenty-eight point three square millimeters eighteen point eight millimeters; one hundred thirteen point zero square millimeters nine point four millimeters; one hundred thirteen point zero square millimeters

OpenStudy (anonymous):

this is really hard! i have like 10 - 20 questions that i am having SOOOO trouble on

jhonyy9 (jhonyy9):

diameter =6 mm so than radius =3 mm circumference =2pir =2pi*3=6pi mm area = pir^2 = pi*3^2 = 9pi mm^2

OpenStudy (anonymous):

thank you so much!

jhonyy9 (jhonyy9):

yw good luck bye

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

would u be able to help me with some more? please :)

jhonyy9 (jhonyy9):

sorry

OpenStudy (anonymous):

so what is it